# Optimal Treatment Allocation Strategies for A/B Testing in Partially Observable Time Series Experiments

### v2

@@ 3.2 The Controlled ARMA Model, Connection to POMDPs, and Estimation of ATE
Controlled ARMA($p,q$). We first introduce the controlled ARMA model, a sub-class of POMDPs, designed to capture carryover effects and partial observability in time series experiments; see Figure 3 for a graphical visualization. The one-dimensional controlled ARMA($p,q$) model is formulated as:
@@
  • The second term in orange on the RHS of model (3.2) incorporates the allocated treatment into the model, affecting the observation $Y_{t}$ at each time. Its treatment effect is measured by the parameter $b$.
  • The last term in purple represents the residual, denoted by $Z_{t}$, which is modeled by a moving average process with the parameters $\{\theta_{j}\}_{j=1}^{q}$, i.e., $Z_{t}=\sum_{j=0}^{q}\theta_{j}\epsilon_{t-j}$. We assume $\theta_{0}=1$, and the white noises $\{\epsilon_{t}\}_{t}$ are i.i.d. with zero mean and variance $\sigma^{2}$.
@@
Finally, different sets of parameters play distinct roles in A/B testing: the autoregressive coefficients ($\{a_{j}\}_{j=1}^{p}$) and the control parameter ($b$) determine the ATE, whereas the moving average coefficients ($\{\theta_{j}\}_{j=1}^{q}$) influence the residual correlation, which in turn determines the optimal design. Formal statements can be found in Lemma 1 and Theorem 1.
@@
In this model: (i) Both the observation $\bm{Y}_{t}$ and the treatment $\bm{U}_{t}$ can be multi-dimensional. (ii) $\bm{X}_{t}$ denotes a vector-valued latent state such that any dependence between the past and future will “funnel” through this latent state. (iii) $\bm{V}_{t}$ and $\bm{W}_{t}$ are the measurement errors. $F,B,H,$ and $C$ are the parameter matrixes, respectively. This model can also be viewed as a variant of the linear state space or dynamic linear model, while incorporating an extra treatment variable $\bm{U}_{t}$.
@@
According to the Wold decomposition theorem (Wold,, 1938), any stationary process can be decomposed into two mutually uncorrelated processes: a linear combination of lags of a white noise process (MA($\infty$) process) and a linear combination of its past values (AR($\infty$) process). The stationarity assumption can typically be satisfied in practice by applying periodic filtering to remove seasonal effects, as detailed in our data analysis in Section 5.2. This underlying principle in time series theory indicates that our model is broadly applicable and can represent a diverse range of linear POMDPs.
@@
We make two remarks. First, Assumption 1 guarantees the ergodicity of the proposed controlled ARMA model, which in turn validates the limits in the definition of the ATE (see (3.1)). Second, as commented earlier, the ATE is exclusively determined by the autoregressive coefficients and the control parameter, and it remains independent of the moving average coefficients. This motivates us to apply the method of moments (e.g., the Yule-Walker method, Yule,, 1927; Walker,, 1931) to estimate the ATE.
@@
We next replace the expectations in (3.4) and (3.5) by their sample moments from $t=p+q+1$ to $T$ and construct $p+2$ estimating equations. Subsequently, we obtain the Yule-Walker estimators $\{\widehat{a}_{j}\}_{j}$ and $\widehat{b}$ for $\{a_{j}\}_{j}$ and $b$, respectively, by which we construct the following estimator for ATE:
@@
By definition, the asymptotic property of (3.6) depends on those of $\{\widehat{a}_{j}\}_{j=1}^{p}$ and $\widehat{b}$. However, deriving their asymptotic variances is challenging, and no closed-form expressions are available to the best of our knowledge. To establish the ATE estimator’s asymptotic MSE, we introduce a small signal asymptotic framework, which will be detailed in the next section.
@@ 3.3 Small Signal Asymptotics, MSEs of ATE Estimators, and Efficiency Indicators
  • Large sample. The first condition is the conventional large sample requirement, which requires the sample size $T$ to grow to infinity. In our ride-sharing example, most experiments last for two weeks, divided into 30-minute or 1-hour intervals, resulting in $T=672$ or $336$ time units.
@@
Under the first large sample condition, the third term in (3.7) – a high-order reminder term – becomes negligible. As such, the first two terms, which measure the discrepancies between the Yule-Walker estimators and their oracle values, become the leading terms. However, deriving their asymptotic variances remains extremely challenging under partial observability.
@@
According to Corollary 1, the asymptotic MSE of the ATE estimator is determined by three factors: (i) the variance of the white noise; (ii) the autoregressive coefficients; and (iii) $\xi_{\pi}$, which measures the percentage of time the new treatment is applied. Different designs affect the ATE’s asymptotic variance only through $\xi_{\pi}$. In other words, designs with the same $\xi_{\pi}$ achieve the same statistical efficiency in estimating the ATE. This uniformity is due to the uncorrelated residuals in the AR model and the small signal condition. Additionally, it turns out that any (asymptotically) balanced design with $\xi_{\pi}=0$ is optimal. This principle holds even when $q>0$, as detailed in Theorem 3 in Section 4. These observations align with the findings of Xiong et al., (2023), highlighting the importance of balancing periodicity in switchback designs under a different model setup.
@@
Additionally, Corollary 2 motivates us to define $\textrm{EI}_{\textrm{AD}}=\sum_{k=1}^{q}\sum_{j=k}^{q}\theta_{j}\theta_{j-k}$ and $\textrm{EI}_{\textrm{AT}}=\sum_{k=1}^{q}(-1)^{k}\sum_{j=k}^{q}\theta_{j}\theta_{j-k}$ as two efficiency indicators. By (3.8), it is immediate to see that
@@
These indicators are useful for comparing the three designs. In practice, one can estimate the moving average coefficients from historical data and plug these estimators into the indicators to determine the most effective design. Alternatively, this procedure can be applied sequentially: use current experimental data up to a specific day to learn the model and estimate the efficiency indicators. Then, apply the optimal design for the following day and continue this cycle by incorporating data from the subsequent day.
3.4 Multivariate Extensions
Next, we extend the univariate controlled ARMA model to its multivariate version, derive asymptotic MSEs of the estimated ATEs, and propose multivariate efficiency indicators.
Controlled VARMA($p,q$). We define the controlled VARMA($p,q$) model as:
$\mathbf{Y}_{t}=\bm{\mu}+\sum_{j=1}^{p}\mathbf{A}_{j}\mathbf{Y}_{t-j}+\mathbf{b}U_{t}+\mathbf{Z}_{t}\,\,\,\,\hbox{and}\,\,\,\,\mathbf{Z}_{t}=\sum_{j=0}^{q}\mathbf{M}_{j}\bm{\epsilon}_{t-j},$ (3.9)
where the bold vectors $\bm{\mu},\mathbf{Y}_{t}$, $\mathbf{Z}_{t}$ and $\bm{\epsilon}_{t}$ denote the $d$-dimensional intercept, observation, residual, and the white noise, respectively. Denote $\bm{\Sigma}$ as the covariance matrix of $\bm{\epsilon}_{t}$. The treatment $U_{t}$ remains binary, taking values in $\{-1,1\}$. Similar to the univariate controlled ARMA model in Section 3.2, model (3.9) contains three sets of parameters: (i) the autoregressive coefficient matrices $\mathbf{A}_{1},\ldots,\mathbf{A}_{p}\in\mathbb{R}^{d\times d}$; (ii) the control coefficient vector $\mathbf{b}\in\mathbb{R}^{d}$; and (iii) the moving average coefficient matrices $\mathbf{M}_{1},\ldots,\mathbf{M}_{q}\in\mathbb{R}^{d\times d}$ and $\mathbf{M}_{0}=\mathbb{I}\in\mathbb{R}^{d\times d}$ as an identity matrix. We next introduce the no unit root assumption for the VARMA model and derive the closed-form expression for the ATE under different treatment allocation strategies.
@@ 3.4 Multivariate Extensions
$\lim_{\begin{subarray}{c}T\rightarrow+\infty\\ \textrm{ATE}\to 0\end{subarray}}\text{MSE}(\sqrt{T}\widehat{\text{ATE}}(\pi_{\text{AT}}))$ $=4\bm{e}^{\top}(\mathbb{I}-\mathbf{A})^{-1}\left(\sum_{j_{1}=0}^{q}\sum_{j_{2}=0}^{q}(-1)^{|j_{2}-j_{1}|}\mathbf{M}_{j_{1}}\bm{\Sigma}\mathbf{M}_{j_{2}}\right)(\mathbb{I}-\mathbf{A})^{-1}\bm{e}.$
@@
$\textrm{EI}_{\text{AD}}&=\bm{e}^{\top}(\mathbb{I}-\mathbf{A})^{-1}\sum_{k=1}^{q}\sum_{j=k}^{q}\mathbf{M}_{j}\bm{\Sigma}\mathbf{M}_{j-k}(\mathbb{I}-\mathbf{A})^{-1}\bm{e}~{}~{}~{}~{}\mbox{and}\\ \textrm{EI}_{\text{AT}}&=\bm{e}^{\top}(\mathbb{I}-\mathbf{A})^{-1}\sum_{k=1}^{q}\sum_{j=k}^{q}(-1)^{k}\mathbf{M}_{j}\bm{\Sigma}\mathbf{M}_{j-k}(\mathbb{I}-\mathbf{A})^{-1}\bm{e}.\end{aligned}$ (3.10)
According to Corollary 3, the efficiency indicators (3.10) enable us to compare the statistical efficiency of the three designs in estimating the ATE in the controlled VARMA model.
It is noteworthy that, similar to the univariate case, the ATE in the multivariate setting also depends on the AR coefficients and the control parameter. However, unlike the univariate setting, where the designs’ relative efficiency is solely determined by MA coefficients, the relative efficiency in the multivariate setting also depends on the AR coefficients. This dependence is reflected by the efficiency indicators in (3.10), additionally involving the coefficient matrix $\mathbf{A}$. It arises as it is challenging to disentangle the effects of the AR coefficients on the ratios of MSEs when comparing two designs in the multivariate setting.
@@ 4 Optimal Treatment Allocation Strategies
This section focuses on the optimal observation-agnostic design, where the ATE estimator derived from the experimental data achieves the smallest asymptotic MSE. Identifying the optimal design is computationally intractable. Each observation-agnostic design is determined by a sequence of treatment allocation strategies $\pi=\{\pi_{t}\}_{t=1}^{T}$, where each $\pi_{t}$ specifies the conditional distribution of $U_{t}$ given $U_{1},\ldots,U_{t-1}$. Consider the class of deterministic treatment allocation strategies where each $\pi_{t}$ is a degenerate distribution. Since $U_{t}$s are binary, there are $2^{t}$ possible $\pi_{t}$ at each time point. Optimizing over such an exponentially growing number of strategies makes the problem NP-hard.
@@
4.1 A Constrained Optimization Approach to Learning Optimal Markov Policies
@@
By definition, both AT and UR are induced by policies within this class. Specifically, setting $\alpha=0$ and $\beta=1$ results in the AT design, whereas $\alpha=\beta=1/2$ yields the UR design. The sequence $\{U_{t}\}_{t}$ forms a Markov chain with binary states. With some calculations, it can be shown that $\xi_{\pi}=(\alpha+\beta-1)/(\beta+1-\alpha)$ in general. To obtain a balanced design, we set $\beta=1-\alpha$, leading to $\xi_{\pi}=0$. When $\alpha=1$ and $\beta=0$, we alternate the initial treatment on a daily basis, yielding the AD design. This indicates the generality of the considered Markov and stationary policy class, which unifies the AD, UR, and AT designs. It remains to identify the optimal $\alpha$ to minimize the asymptotic MSE of resulting the ATE estimator, which — under the small signal asymptotic framework — can be derived as
@@
where $c_{0}={\sum_{j=0}^{q}\theta^{2}_{j}}/{(1-a)^{2}}$ and $c_{k}={\sum_{j=k}^{q}\theta_{j}\theta_{j-k}}/{(1-a)^{2}}$ under the controlled ARMA($p,q$) model, while under the controlled VARMA($p,q$) model we have $c_{0}=\bm{e}^{\top}(\mathbb{I}-\mathbf{A})^{-1}\left(\sum_{j=0}^{q}\mathbf{M}_{j}\Sigma\mathbf{M}_{j}\right)(\mathbb{I}-\mathbf{A})^{-1}\bm{e}$ and $c_{k}=\bm{e}^{\top}(\mathbb{I}-\mathbf{A})^{-1}\left(\sum_{j=k}^{q}\mathbf{M}_{j}\Sigma\mathbf{M}_{j-k}\right)(\mathbb{I}-\mathbf{A})^{-1}\bm{e}$. See Appendix E for details. This asymptotic MSE form motivates us to compute $\alpha$ by solving the following constrained $q$-order polynomial optimization:
@@
The above optimization can be efficiently solved using existing convex optimization techniques, such as the limited-memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS) algorithm (Liu and Nocedal,, 1989). Notice that $c_{k}$ and the optimal number of AR and MA lags, $p$ and $q$, depend on the true model, which are typically unknown. However, as discussed in Section 3.3, they can be effectively estimated or evaluated using historical or initial experimental data in practice. The optimal $p$ and $q$ can be selected based on the Akaike information criterion (AIC, Akaike,, 1974) or the Bayesian information criterion (BIC, Schwarz,, 1978).
4.2 A Reinforcement Learning Approach to Learning Optimal $q$-dependent Policies
In this section, we consider the more general history-dependent class of treatment allocation strategies beyond Markov policies and propose an RL algorithm to identify the optimal treatment allocation strategy, denoted by $\pi^{*}$. The primary objective of RL is to learn an optimal policy, a mapping from time-varying environmental features (referred to as state) to decision rules about which treatment to administer (referred to as action), in order to maximize the expected cumulative outcome (where each intermediate outcome is referred to as a reward). Most existing RL algorithms estimate the optimal policy by modeling these state-action-reward triplets over time as an MDP, wherein each reward and future state are independent of the past history given the current state-action pair.

### v4

@@ 3.2 The Controlled ARMA Model, Connection to POMDPs, and Estimation of ATE
Controlled ARMA($p,q$). We first introduce the controlled ARMA model, a sub-class of POMDPs, designed to capture carryover effects and partial observability in online experiments; see Figure 3 for a graphical visualization. The one-dimensional controlled ARMA($p,q$) model is formulated as:
@@
  • The second term in orange on the RHS of model (3.2) incorporates the treatment into the model, affecting the observation $Y_{t}$ at each time. Its treatment effect is measured by the parameter $b$.
  • The last term in purple represents the residual, denoted by $Z_{t}$, which is modeled by a moving average process with the parameters $\{\theta_{j}\}_{j=1}^{q}$, i.e., $Z_{t}=\sum_{j=0}^{q}\theta_{j}\epsilon_{t-j}$. We assume the white noises $\{\epsilon_{t}\}_{t}$ are i.i.d. with zero mean and variance $\sigma^{2}$.
@@
Finally, different sets of parameters play different roles in A/B testing: the autoregressive coefficients ($\{a_{j}\}_{j=1}^{p}$) and the control parameter ($b$) determine the ATE, whereas the moving average coefficients ($\{\theta_{j}\}_{j=1}^{q}$) influence the residual correlation, which in turn determines the optimal design. Formal statements can be found in Lemma 1 and Theorem 1.
@@
In this model: (i) Both the observation $\bm{Y}_{t}$ and the treatment $\bm{U}_{t}$ can be multi-dimensional. (ii) $\bm{X}_{t}$ denotes a vector-valued latent state such that any dependence between the past and future will “funnel” through this latent state. (iii) $\bm{V}_{t}$ and $\bm{W}_{t}$ are the measurement errors. $F,B,H,$ and $C$ are the parameter matrices, respectively. This model can also be viewed as a variant of the linear state space or dynamic linear model, which incorporates an extra treatment variable $\bm{U}_{t}$.
@@
According to the Wold decomposition theorem (Wold,, 1938), any stationary process can be decomposed into two mutually uncorrelated processes: a linear combination of lags of a white noise process (MA($\infty$) process) and a linear combination of its past values (AR($\infty$) process). The stationarity assumption can typically be satisfied in practice by applying periodic filtering to remove seasonal effects, as detailed in Section 3.4 and our data analysis in Section 5.3. This underlying principle in time series theory indicates that our model is broadly applicable and can represent a diverse range of linear POMDPs.
@@
We make three remarks. First, Assumption 1 guarantees the ergodicity of the proposed controlled ARMA model, which in turn validates the limits in the definition of the ATE (see (3.1)). Second, the ATE can be decomposed into the sum of $2b+2ab/(1-a)$ where the first term corresponds to the direct effect of $U_{t}$ on $Y_{t}$ and the second term represents the indirect effect mediated by the past observations $\{Y_{t-j}\}_{j\geq 1}$. Third, as commented earlier, the ATE is exclusively determined by the autoregressive coefficients and the control parameter, and it remains independent of the moving average coefficients. This motivates us to apply the method of moments (e.g., the Yule-Walker method, Yule,, 1927; Walker,, 1931) to estimate the ATE.
@@
We next replace the expectations in (3.4) and (3.5) by their sample moments from $t=p+q+1$ to $T$ and construct $p+2$ estimating equations. Subsequently, we solve these equations to obtain the Yule-Walker estimators $\{\widehat{a}_{j}\}_{j}$ and $\widehat{b}$ for $\{a_{j}\}_{j}$ and $b$, respectively, by which we construct the following estimator for ATE:
@@
By definition, the asymptotic property of (3.6) depends on those of $\{\widehat{a}_{j}\}_{j=1}^{p}$ and $\widehat{b}$. However, deriving their asymptotic variances is extremely challenging, and no closed-form expressions are available to the best of our knowledge. To establish the ATE estimator’s asymptotic MSE, we introduce a small signal asymptotic framework, detailed in the next section.
@@ 3.3 Small Signal Asymptotics, MSEs of ATE Estimators, and Efficiency Indicators
  • Large sample. The first condition is the conventional large sample condition, which requires the sample size $T$ to grow to infinity. In our ride-sharing example, most experiments last for two weeks, divided into 30-minute or 1-hour intervals, resulting in $T=672$ or $336$ time units.
@@
Under the first large sample condition, the third term in (3.7) – a high-order reminder term – becomes negligible. As such, the first two terms, which measure the discrepancies between the Yule-Walker estimators and their oracle values, become the leading terms. However, as mentioned earlier, deriving their asymptotic variances remains extremely challenging under partial observability.
@@
According to Corollary 1, the asymptotic MSE of the ATE estimator is determined by three factors: (i) the variance of the white noise; (ii) the autoregressive coefficients; and (iii) $\xi_{\pi}$, which measures the percentage of time the new treatment is applied. Different designs affect the ATE’s asymptotic variance only through $\xi_{\pi}$. In other words, designs with the same $\xi_{\pi}$ achieve the same statistical efficiency in estimating the ATE. This uniformity is due to the uncorrelated residuals in the AR model. Additionally, it turns out that any (asymptotically) balanced design with $\xi_{\pi}=0$ is optimal. This principle holds even when $q>0$, as detailed in Theorem 3 in Section 4. These observations align with the findings of Xiong et al., (2023), highlighting the importance of balancing periodicity in switchback designs under a different model setup.
@@
Finally, Corollary 2 motivates us to define two efficiency indicators $\textrm{EI}_{\textrm{AD}}=\sum_{k=1}^{q}\sum_{j=k}^{q}\theta_{j}\theta_{j-k}$ and $\textrm{EI}_{\textrm{AT}}=\sum_{k=1}^{q}(-1)^{k}\sum_{j=k}^{q}\theta_{j}\theta_{j-k}$. By (3.8), it is immediate to see that
@@
These indicators are useful for comparing the three designs. In practice, one can estimate the moving average coefficients from historical or initial experimental data and plug these estimators into the indicators to determine the most effective design among the three. In Section 4, we discuss methodologies to search the optimal design within the broader class of observation-agnostic designs, beyond just these three.
To conclude this section, we note that although our asymptotic derivation and the subsequently proposed designs rely on the small signal condition, our proposal remains effective even with large treatment effects. This is because in these scenarios, the design problem itself might not be that critical, and any reasonable design should be able to detect these effects. Therefore, our proposal remains a safe option to use, regardless of whether this assumption holds or not.
3.4 Extensions
In this section, we extend the univariate controlled ARMA model by accommodating multivariate observations and exogenous variables, derive asymptotic MSEs of the estimated ATEs, and propose the resulting efficiency indicators.
Controlled VARMA($p,q$). We define the controlled VARMA($p,q$) model with an additional exogenous variable $\bm{E}_{t}$ as:
$\mathbf{Y}_{t}=\bm{\mu}+\sum_{j=1}^{p}\mathbf{A}_{j}\mathbf{Y}_{t-j}+\mathbf{b}U_{t}+\bm{C}\bm{E}_{t}+\mathbf{Z}_{t}\,\,\,\,\hbox{and}\,\,\,\,\mathbf{Z}_{t}=\sum_{j=0}^{q}\mathbf{M}_{j}\bm{\epsilon}_{t-j},$ (3.9)
where the bold vectors $\bm{\mu},\mathbf{Y}_{t}$, $\mathbf{Z}_{t}$ and $\bm{\epsilon}_{t}$ denote the $d$-dimensional intercept, observation, residual and the white noise, respectively. The treatment $U_{t}$ remains binary, taking values in $\{-1,1\}$. The purpose of introducing the extra exogenous variable $\bm{E}_{t}$ is to further enhance model flexibility. This variable remains unaffected by the treatments and can be regarded as the “non-stationary” components of the model, accounting for a broad range of temporal factors, such as the daily seasonal trends (see Section 5.3 for the construction of $\bm{E}_{t}$).
Model (3.9) contains four sets of parameters: (i) the autoregressive coefficient matrices $\mathbf{A}_{1},\ldots,\mathbf{A}_{p}\in\mathbb{R}^{d\times d}$; (ii) the control coefficient vector $\mathbf{b}\in\mathbb{R}^{d}$; (iii) the moving average coefficient matrices $\mathbf{M}_{1},\ldots,\mathbf{M}_{q}\in\mathbb{R}^{d\times d}$ and $\mathbf{M}_{0}=\mathbb{I}\in\mathbb{R}^{d\times d}$ as an identity matrix; (iv) the coefficient matrix $\bm{C}$ for the extra exogenous variable.
We next introduce the no unit root assumption for the VARMA model and derive the closed-form expression for the ATE using different treatment allocation strategies.
@@ 3.4 Extensions
$\lim_{\begin{subarray}{c}T\rightarrow+\infty\\ \textrm{ATE}\to 0\end{subarray}}\text{MSE}(\sqrt{T}\widehat{\text{ATE}}(\pi_{\text{AT}}))$ $=4\bm{e}^{\top}(\mathbb{I}-\mathbf{A})^{-1}\left(\sum_{j_{1}=0}^{q}\sum_{j_{2}=0}^{q}(-1)^{|j_{2}-j_{1}|}\mathbf{M}_{j_{1}}\bm{\Sigma}\mathbf{M}_{j_{2}}\right)(\mathbb{I}-\mathbf{A})^{-1}\bm{e},$
where $\bm{\Sigma}$ denotes the covariance matrix of $\bm{\epsilon}_{t}$.
@@
$\textrm{EI}_{\text{AD}}&=\bm{e}^{\top}(\mathbb{I}-\mathbf{A})^{-1}\sum_{k=1}^{q}\sum_{j=k}^{q}\mathbf{M}_{j}\bm{\Sigma}\mathbf{M}_{j-k}(\mathbb{I}-\mathbf{A})^{-1}\bm{e}~{}~{}~{}~{}\mbox{and}\\ \textrm{EI}_{\text{AT}}&=\bm{e}^{\top}(\mathbb{I}-\mathbf{A})^{-1}\sum_{k=1}^{q}\sum_{j=k}^{q}(-1)^{k}\mathbf{M}_{j}\bm{\Sigma}\mathbf{M}_{j-k}(\mathbb{I}-\mathbf{A})^{-1}\bm{e}.\end{aligned}$
According to Corollary 3, they enable us to compare the statistical efficiency of the three designs in estimating the ATE in the controlled VARMA model.
@@ 4 Optimal Treatment Allocation Strategies
This section focuses on the optimal observation-agnostic design, where the ATE estimator derived from the experimental data achieves the smallest asymptotic MSE. Identifying the optimal design is computationally intractable. To elaborate, each observation-agnostic design is determined by a sequence of treatment allocation strategies $\pi=\{\pi_{t}\}_{t=1}^{T}$, where each $\pi_{t}$ specifies the conditional distribution of $U_{t}$ given $U_{1},\ldots,U_{t-1}$. Consider the class of deterministic treatment allocation strategies where each $\pi_{t}$ is a degenerate distribution. Since $U_{t}$s are binary, there are $2^{t}$ possible $\pi_{t}$ at each time point. Optimizing over such an exponentially growing number of strategies makes the problem NP-hard.
@@
4.1 A Constrained Optimization Approach
@@
By definition, both AT and UR are induced by policies within this class. Specifically, setting $\alpha=0$ and $\beta=1$ results in the AT design, whereas $\alpha=\beta=1/2$ yields the UR design. When $\alpha=1$, $\beta=0$, and we alternate the initial treatment on a daily basis, it yields the AD design. This indicates the generality of the considered Markov and stationary policy class, which unifies the AD, UR and AT designs.
Additionally, the sequence $\{U_{t}\}_{t}$ forms a Markov chain with binary states. With some calculations, it can be shown that $\xi_{\pi}=(\alpha+\beta-1)/(\beta+1-\alpha)$ in general. To obtain a balanced design, we set $\beta=1-\alpha$, leading to $\xi_{\pi}=0$. It remains to identify the optimal $\alpha$ to minimize the asymptotic MSE of resulting the ATE estimator, which — under the small signal asymptotic framework — can be derived as
@@
where $c_{0}={\sum_{j=0}^{q}\theta^{2}_{j}}/{(1-a)^{2}}$ and $c_{k}={\sum_{j=k}^{q}\theta_{j}\theta_{j-k}}/{(1-a)^{2}}$ under the controlled ARMA($p,q$) model, while under the controlled VARMA($p,q$) model we have
$c_{0}$ $=\bm{e}^{\top}(\mathbb{I}-\mathbf{A})^{-1}\left(\sum_{j=0}^{q}\mathbf{M}_{j}\Sigma\mathbf{M}_{j}\right)(\mathbb{I}-\mathbf{A})^{-1}\bm{e},$ (4.2)
$c_{k}$ $=\bm{e}^{\top}(\mathbb{I}-\mathbf{A})^{-1}\left(\sum_{j=k}^{q}\mathbf{M}_{j}\Sigma\mathbf{M}_{j-k}\right)(\mathbb{I}-\mathbf{A})^{-1}\bm{e}.$
See Appendix E for more details. This asymptotic MSE formula motivates us to compute $\alpha$ by solving the following constrained $q$-order polynomial optimization:
@@
The above optimization can be efficiently solved using existing convex optimization techniques, such as the limited-memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS) algorithm (Liu and Nocedal,, 1989). Notice that $c_{k}$ and the optimal number of AR and MA lags, $p$ and $q$, depend on the true model, which are typically unknown. However, as discussed in Section 3.3, they can be effectively estimated or evaluated using historical data in practice. For instance, the optimal $p$ and $q$ can be selected based on the Akaike information criterion (AIC, Akaike,, 1974) or the Bayesian information criterion (BIC, Schwarz,, 1978). Alternatively, this procedure can be applied sequentially: use current experimental data up to a specific day to learn $\{c_{k}\}_{k}$, $p$ and $q$ to estimate the optimal design. Then, this design will be applied on the following day, and the estimating procedure will continue by incorporating data from the subsequent day.
4.2 A Reinforcement Learning Approach
In this section, we consider the more general history-dependent policy class and propose an RL algorithm to identify the unrestricted optimal treatment allocation strategy $\pi^{*}$. The primary objective of RL is to learn an optimal policy, a mapping from time-varying environmental features (referred to as state) to decision rules about which treatment to administer (referred to as action), in order to maximize the expected cumulative outcome (where each intermediate outcome is referred to as a reward). Most existing RL algorithms estimate the optimal policy by modeling these state-action-reward triplets over time as an MDP, wherein each reward and future state are independent of the past history given the current state-action pair.